40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8501     Entry time: Sat Apr 27 00:29:40 2013
Author: Koji 
Type: Update 
Category: LSC 
Subject: LSCoffset script fixed 

Prior to the locking trials...

scripts/LSC/LSCoffset script had behaved peculiarly:

This script spawns LSC/offset3 in order to remove the dark offset from the channels.
How ever the offsets had been nulled every other PDs
(i.e. The offsets REFL11 I&Q were nulled.
The offsets REFL33 I&Q had been left untouched
The offset REFL55 I&Q had been nulled
and so on.)

I found that the script run many instances of "offset3" scripts in background.
It seemed that tdsavg did not like too many averaging channels at once.

So the "&"s in the LSCoffsets were removed and now the script runs much more slowly,
but works for all of the PDs listed.

I think I have never seen the offsets in REFL33 and REFL165 nulled down to this level before.

ELOG V3.1.3-